OpenStudy (anonymous):

solve 2x-7=6x-15 a 2 b 5 c 3 d 6

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

x = 2

OpenStudy (adilalvi):

a

OpenStudy (anonymous):

Just take like terms together, and do your thing ;)

OpenStudy (texaschic101):

2x - 7 = 6x - 15 -- add 7 to both sides 2x = 6x - 15 + 7 -- subtract 6x from both sides 2x - 6x = - 15 + 7 -- combine like terms -4x = - 8 -- divide both sides by -4 x = 2 I see you already have the answer....I just wanted to show you how to do it :)

OpenStudy (anonymous):

Nice one @texaschic101! (: